# The Mind in the Machine: A Survey of Incorporating Psychological Theories in LLMs

### v1

@@ 5.1.3 Memory and Cognitive Evaluation
Assessing memory and cognition is crucial given LLMs’ limited capacity and risk of catastrophic forgetting. Cognitive load is measured by Xu et al. (2024b) on jail-breaking and Zeng et al. (2024b) on memorization patterns. Memory is measured by Li et al. (2023) on parametric knowledge, by Zhang et al. (2024a) on n-back tasks and by Timkey & Linzen (2023) on capacity. Meanwhile, cognitive development and reasoning capabilities have been assessed through cognitive maturity (Laverghetta Jr. & Licato, 2022), subjective similarity (Malloy et al., 2024), reasoning strategies (Mondorf & Plank, 2024; Yuan et al., 2023), decision-making (Ying et al., 2024), and ToM (Jung et al., 2024).
@@ 5.1.4 Personality Capability
Personality consistency examines how stably LLMs maintain traits across contexts. Frisch & Giulianelli (2024) show LLMs with asymmetric profiles vary in Big Five traits – Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism, while Amidei et al. (2025) find language switching alters GPT-4o’s traits on the Eysenck Personality Questionnaire Revised, underscoring challenges in preserving stable traits and reducing context dependence. Parallel research examines how LLMs display and control personality traits. Jiang et al. (2024) show LLMs can express distinct Big Five traits recognized by human evaluators. Mao et al. (2024) introduce PersonalityEdit, revealing difficulties in maintaining consistent alignment for Neuroticism, Extraversion, and Agreeableness, while Hu & Collier (2024) find persona-based prompting improves annotation accuracy.
@@ 5.1.5 Bias and Ethics Evaluation
Evaluating biases and ethical risks is crucial for responsible AI that avoids reinforcing harmful social patterns. As LLMs increasingly shape public discourse, thorough assessments are essential to prevent discriminatory outputs and promote equitable benefits across diverse communities. Recent work tests LLMs on gender (Oba et al., 2024; Zhao et al., 2024), broader social biases (Shin et al., 2024; Lee et al., 2023; Nozza et al., 2022), toxic content (Gehman et al., 2020; Luong et al., 2024; Hui et al., 2024), and harmful stereotypes (Shrawgi et al., 2024; Huang & Xiong, 2024; Grigoreva et al., 2024), establishing benchmarks across cultures and languages. Evidence also suggests that LLMs replicate social identity biases, mirroring human tendencies toward ingroup favoritism and outgroup hostility (Hu et al., 2025a; Dong et al., 2024a) – patterns central to social identity theory, which posits that group membership shapes self-concept and intergroup behavior (Tajfel, 1979).
@@ 5.2 Task Performance Enhancement
Building on the benchmarks, we review how psychological insights boost LLM task performance. We highlight techniques that enhance complex reasoning and enrich dialogue, illustrating how psychology improves capabilities and alignment across applications.
@@ 5.2.1 Reasoning Enhancement
LLMs often struggle with complex reasoning – social inference (Liu et al., 2024a), logical errors (Turpin et al., 2023; McKenna et al., 2023), hallucinations (Huang et al., 2025; Ai et al., 2024a), and multi-step planning (Wang et al., 2024a). Psychological theories of human cognition offer frameworks to address these issues by implementing analogous cognitive mechanisms. For instance, Dual-process theories, a social cognition framework, distinguish between fast (System 1) and slow (System 2) reasoning (Kahneman, 2011), offering a blueprint for LLM improvement. Chain-of-thought prompting (Wei et al., 2022) operationalizes System 2 via intermediate steps, while DynaThink (Pan et al., 2024) dynamically selects rapid or thorough inference. Tree of Thoughts (Yao et al., 2024) further explores multiple reasoning paths concurrently.
Similarly, Self-reflectionintrospection focused on the self-concept (Phillips, 2020) – has guided LLM enhancements in hallucination mitigation (Ji et al., 2023), translation (Chen et al., 2024a; Wang et al., 2024e), question-answering (Li et al., 2024a; Zhang et al., 2024f; Kassner et al., 2023), and math reasoning (Zhang et al., 2024f). Approaches include iterative self-assessment (Ji et al., 2023; Yan et al., 2024), task decomposition (Wang et al., 2024e; Zhang et al., 2024d), reflection-driven self-training (Dou et al., 2024), and confidence-tuned reward functions (Xu et al., 2024c). Additionally, ToM adaptations boost LLMs’ interpersonal reasoning, aiding missing knowledge inference (Bortoletto et al., 2024), common ground alignment (Qiu et al., 2024), and cognitive modeling (Wu et al., 2024a).
Beyond social reasoning, Perception, Attention, and Memory support coherence and retrieval. Kojima et al. (2022) uses “Let’s think step by step” prompts for top-down reasoning. Maharaj et al. (2023); Yu et al. (2022) leverages selective attention to detect hallucinations and extract relation. Zhu et al. (2024) employs recitation for retrieval, and Park & Bak (2024) introduce separate short- and long-term memory modules. Finally, Wang et al. (2024c); Chi et al. (2023) improve complex reasoning via symbolic and adaptive memory structures.
Another challenge is multi-step reasoning with external knowledge, as in retrieval-augmented generation (RAG). Hippocampal indexing theory (Teyler & DiScenna, 1986), viewing the hippocampus as a pointer to neocortical memory, enhances RAG (Gutierrez et al., 2024) and counterfactual reasoning (Miao et al., 2024a). Meanwhile, self-reflection and meta-cognition (Phillips, 2020; Flavell, 1979), supporting iterative introspection, improve retrieval (Asai et al., 2024) and multi-step inference (Zhou et al., 2024).
@@ 5.2.2 Dialogue Understanding and Generation
Dialogue understanding is a key area where personality psychology aids trait-based inferences from user interactions. NLP research has explored dynamic ways to measure personality beyond structured tests. The Myers–Briggs Type Indicator (MBTI), a self-report questionnaire that makes pseudo-scientific claims to categorize individuals into 16 distinct personality types, remains popular (Rao et al., 2023; Yang et al., 2023), while PsychoGAT (Yang et al., 2024) gamifies MBTI, and PADO (Yeo et al., 2025) adopts a Big Five-based multi-agent approach. Beyond assessments, traits guide dialogue generation: Huang & Hadfi (2024) show higher agreeability improves negotiation, while Cheng et al. (2023) reveal social and racial biases in persona creation, raising representational concerns.
Dialogue generation research further incorporates personality to improve coherence, empathy, and consistency. Pal et al. (2025); Chen et al. (2025a) leveraged Reddit-based journal entries to model Big Five traits in large-scale dialogue datasets. Other efforts improve persona consistency without referencing explicit psychological theory (Wu et al., 2025b; Takayama et al., 2025). Similarly, personality is used to improve truthfulness, consistency, and context-aware generation in LLMs, as further detailed in Appendix C. These approaches support personality alignment but lack grounding in deeper psychological theory.
5.3 Collaborative and Multi-Agent Frameworks
Beyond task-specific capabilities, the surge in multi-agent LLM frameworks reflects a growing emphasis on collaborative decision-making – where modeling social dynamics is crucial. Social and personality psychology theories offer key insights into designing agent interaction, negotiation, and consensus, guiding more socially intelligent LLM systems.
@@ 5.3 Collaborative and Multi-Agent Frameworks
Persuasion models (Petty & Cacioppo, 2012) illustrate how central or peripheral routes shape attitudes in collaborative settings. Leveraging this, Gollapalli & Ng (2025) merges persuasive dialog acts with reinforcement learning, Furumai et al. (2024) combines LLM strategies and retrieval, Qin et al. (2024); Jin et al. (2024a) emphasize credibility-aware generation, and Zeng et al. (2024a) uncovers LLM vulnerabilities. Multi-agent research simulates personality-driven negotiation (Huang & Hadfi, 2024; Hu et al., 2025b), boosts truthfulness via structured debates (Khan et al., 2024), and curates argument-strength datasets (Rescala et al., 2024).
@@
ToM complements social influence by enabling agents to grasp others’ mental states. Some efforts integrate belief tracking (Sclar et al., 2023) and coordination (Wang et al., 2022; Sclar et al., 2022), while recent approaches (Wilf et al., 2024; Jung et al., 2024) refine ToM via task decomposition and recursive simulation (Sarangi et al., 2025).
@@
Recent work on persona-driven LLM agents focuses on simulating diverse perspectives, persona alignment, and socially intelligent interactions. Han et al. (2024) introduce PSYDIAL (Big Five-based Extraversion), Castricato et al. (2025) present PERSONA (1,586 synthetic personas), and Wu et al. (2025a) release the RAIDEN Benchmark (40K multi-turn dialogues). Agents also model opinion dynamics (Wang et al., 2025) and evaluate social intelligence (Chen et al., 2024b), with RoleLLM (Wang et al., 2024b), Character100 (Wang et al., 2024d), and persona-aware graph transformers (Mahajan & Shaikh, 2024) further supporting multi-party simulations. Additionally, Kumarage et al. (2025) simulate multi-turn social engineering attacks with LLM agents of varied personality traits, highlighting how psychological profiles shape user vulnerability and the need for more robust defenses against personalized manipulation.
@@ 6.1 How Does Current LLM Development Harness Psychological Theories?
From the review from Section 2-5, we observe psychological theories have been incorporated into LLM development in stage-specific ways, with uneven coverage across theoretical domains. Figure 1 maps this integration across the development stages.
In early stages, data selection and pretraining, developmental psychology is often referenced. Its emphasis on staged learning and knowledge acquisition aligns with curriculum learning and progressive data exposure, mirroring human developmental trajectories. In post-training, especially RLHF, behavioral psychology ideas are most prominent. Conditioning, reinforcement schedules, and reward design are commonly used to guide model alignment with human preferences. In evaluation and application, theories from social psychology, personality psychology, and psycholinguistics are commonly cited, reflecting a focus on interaction patterns, user modeling, and linguistic variation – areas traditionally explored within these sub-fields. Their prominence in later stages aligns with their emphasis on human-centered communication. Cognitive psychology appears across all stages, particularly in modeling internal mechanisms such as reasoning, memory, and attention. Its breadth makes it a foundational influence, though less visible in agentic interaction settings.
The observed unevenness in integration reflects, perhaps a gap, but more probably a functional alignment – some psychological domains are naturally better suited for certain stages of model development. Meanwhile, these trends expose under-explored opportunities, which motivate the research questions that follow.
@@ 6.2 What Untapped Psychological Insights Could Advance LLM Development?
Although psychological theory is increasingly applied in LLM research, its use remains simplified and uneven. As shown in Tables A, A, and A, many theories are under-utilized despite their potential to improve model behavior and interpretability. Below, we outline theories in four key areas that deserve greater attention in future LLM research.
Social psychology remains underutilized in areas like group dynamics and self and identity, limiting personalization, adaptability, and inclusivity. Prompting LLMs to adopt specific social identities can reduce bias (Dong et al., 2024a) and mirror human-like ingroup favoritism (Hu et al., 2025a). Incorporating social identity frameworks could enhance user alignment in identity-sensitive contexts (Chen et al., 2020). Likewise, while bias detection is common, classic social influence theories (e.g., conformity, obedience) and attitude change theories (e.g., balance theory, cognitive dissonance) are rarely applied to interaction dynamics or bias mitigation, despite their relevance to ethical and socially adaptive behavior. Additionally, malicious actors leveraging social influence can severely undermine trust in digital spaces (Zeng et al., 2024a; Liu et al., 2025; Ai et al., 2024b), highlighting the potential of constructs like inoculation theory to proactively guard against manipulative strategies.
Behavioral psychology inspires RLHF, yet key concepts like partial reinforcement, which improves behavior persistence (Ferster, 1966; Jensen, 1961), and shaping, which supports gradual learning through successive approximations (Love et al., 2009), are overlooked. Current RLHF relies on uniform rewards, yet behavioral theory warns that flawed rewards can lead to reward hacking. Adding reward variability may reduce premature convergence and improve alignment with human intent (Dayan & Daw, 2008; Amodei et al., 2016).
@@
Cognitive psychology remains underused, particularly Schema Theory, which holds that humans store knowledge as dynamic, structured representations formed through repeated experience (Anderson & Pearson, 1984), guiding inference, memory, and learning. Recent work explores schema-inspired methods for compressing user histories and modeling knowledge activation cycles (Panagoulias et al., 2024; Xia et al., 2024), though these remain peripheral. Further integration may improve long-term context handling and generalization.

### v2

@@ 5.1.3 Memory and Cognitive Evaluation
Assessing memory and cognition is crucial given LLMs’ limited capacity and risk of catastrophic forgetting. Memory is measured on parametric knowledge (Li et al., 2023), n-back tasks (Zhang et al., 2024a), capacity (Timkey and Linzen, 2023) and cognitive load (Fu et al., 2025; Xu et al., 2024b; Zeng et al., 2024b). Meanwhile, cognitive development is assessed through cognitive maturity (Laverghetta Jr. and Licato, 2022; Wang et al., 2025b), word acquisition (Chang and Bergen, 2022), subjective similarity (Malloy et al., 2024), reasoning strategies (Mondorf and Plank, 2024; Yuan et al., 2023; Ying et al., 2024), zone of proximal development (Cui and Sachan, 2025) and perception (Jung et al., 2024; Zhou et al., 2025b).
@@ 5.1.4 Personality Capability
Personality consistency examines how stably LLMs maintain traits across contexts. Frisch and Giulianelli (2024) show LLMs with asymmetric profiles vary in Big Five traits, while Amidei et al. (2025) find language switching alters GPT-4o’s Eysenck Personality Questionnaire Revised traits, underscoring challenges in perserving consistency. Parallel research examines how LLMs display and control personality traits. Jiang et al. (2024) show LLMs express distinct traits labeled by human evaluators. Mao et al. (2024) reveals difficulties in alignment for Neuroticism, Extraversion and Agreeableness. Han et al. (2025); Lee et al. (2025b); Li et al. (2025c, d) assess and improve consistency through alignment with psychometrical training data, while Hu and Collier (2024) find persona-based prompting improves annotation accuracy.
@@ 5.1.5 Bias and Ethics Evaluation
Evaluating biases and ethical risks is crucial for responsible AI that avoids reinforcing harmful social patterns. As LLMs increasingly shape public discourse, thorough assessments are essential to prevent discriminatory outputs and promote equitable benefits across diverse communities. Recent work tests LLMs on gender (Oba et al., 2024; Zhao et al., 2024), broader social biases (Abolghasemi et al., 2025; Huang et al., 2025a; Lee et al., 2023; Muti et al., 2025; Nozza et al., 2022; Shin et al., 2024; Zhao et al., 2025), toxic content (Gehman et al., 2020; Huang et al., 2025c; Hui et al., 2024a; Luong et al., 2024), and harmful stereotypes (Grigoreva et al., 2024; Huang and Xiong, 2024; Hui et al., 2024b; Shrawgi et al., 2024), establishing benchmarks across cultures and languages. Evidence also suggests that LLMs replicate social identity biases, mirroring human tendencies toward ingroup favoritism and outgroup hostility (Borah et al., 2025; Dong et al., 2024a; Hu et al., 2025b; Wang et al., 2025c) – patterns central to social identity theory, which posits that group membership shapes self-concept and intergroup behavior (Tajfel, 1979).
@@ 5.2 Task Performance Enhancement
Building on the benchmarks, we review how psychological insights are used to improve LLMs performance on complex reasoning and enrich dialogue, which illustrate how psychology improves capabilities and alignment across applications.
@@ 5.2.1 Reasoning Enhancement
LLMs often struggle with complex reasoning: social inference (Liu et al., 2024a), logical errors (McKenna et al., 2023; Turpin et al., 2023), hallucinations (Ai et al., 2024a; Huang et al., 2025b), and multi-step planning (Wang et al., 2024a). Researchers address these issues by implementing analogous cognitive mechanisms. For instance, Dual-process theory, a social cognition framework, distinguishes between fast (System 1) and slow (System 2) reasoning (Kahneman, 2011), offers a blueprint for LLM improvement. Chain-of-thought prompting (Wei et al., 2022) operationalizes System 2 via intermediate steps, while DynaThink (Pan et al., 2024) dynamically selects rapid or thorough inference. Tree of Thoughts (Yao et al., 2024) further explores multiple reasoning paths concurrently. Yang et al. (2025) combine separate verifier as System 2. More recent applications includes hallucination mitigation (Cheng et al., 2025c), real-time human-AI collaboration (Zhang et al., 2025b), multi-hop QA (Cheng et al., 2025a), emotion consistency (Wei et al., 2025), decoder-level LLMs merging (Hu et al., 2025a) and cost-efficient RL framework (Cheng et al., 2025b).
Similarly, Self-reflection and Meta-cognition, introspection focused on the self-concept (Flavell, 1979; Phillips, 2020), has guided LLM enhancements in hallucination mitigation (Ji et al., 2023; Lu et al., 2025), translation (Chen et al., 2024a; Wang et al., 2024e), tool use (Li et al., 2025b), question-answering (Kassner et al., 2023; Li et al., 2024a; Zhang et al., 2024f), retrieval-augmented-generation(RAG) (Asai et al., 2024; Zhou et al., 2024) and math reasoning (Zhang et al., 2024f). Approaches include iterative self-assessment (Ji et al., 2023; Wu et al., 2025c; Yan et al., 2024), task decomposition (Wang et al., 2024e; Zhang et al., 2024d), self-training (Dou et al., 2024), in context learning (Ji-An et al., 2025), and confidence-tuned reward functions (Xu et al., 2024c). Moreover, ToM adaptations boost LLMs’ interpersonal reasoning, aiding missing knowledge (Bortoletto et al., 2024), common ground alignment (Qiu et al., 2024), and cognitive modeling (Wu et al., 2024a).
Beyond social reasoning, perception, attention, and memory support coherence and retrieval. Kojima et al. (2022) uses “think step by step” prompts for top-down reasoning. Li et al. (2025a) employed Piaget’s constructivist theory on a systematic memory system for reading comprehension. Chen et al. (2025a); Maharaj et al. (2023); Yu et al. (2022) leverages selective attention and working memory to detect hallucinations and extract relation. Zhu et al. (2024) employs recitation for retrieval, and Park and Bak (2024) introduce short/long-term memory modules. Chi et al. (2023); Diao et al. (2025); Wang et al. (2024c) improve reasoning via symbolic, adaptive and working memory structures. Lastly, hippocampal indexing theory (Teyler and DiScenna, 1986), viewing the hippocampus as a pointer to neocortical memory, informs multi-step reasoning with external knowledge (Fang et al., 2025; Gutierrez et al., 2024) and counterfactual reasoning (Miao et al., 2024a).
@@ 5.2.2 Dialogue Understanding and Generation
In dialogue understanding, personality psychology aids trait-based inferences from user interactions. NLP research has explored dynamic ways to measure personality beyond structured tests. The Myers–Briggs Type Indicator (MBTI), a self-report questionnaire that makes pseudo-scientific claims to categorize individuals into 16 distinct personality types, remains popular (Rao et al., 2023; Yang et al., 2023), while PsychoGAT (Yang et al., 2024) gamifies MBTI, and PADO (Yeo et al., 2025) adopts a Big Five-based multi-agent approach. Beyond assessments, traits guide dialogue generation: Huang and Hadfi (2024) show higher agreeability improves negotiation, while Cheng et al. (2023) reveal social and racial biases in persona creation, raising representational concerns.
Dialogue generation research further incorporates personality to improve coherence, empathy, and consistency. Dan et al. (2025) enforced trait-consistent generation via Big Five trait specialization loss. Joshi et al. (2025) augmented personas with rationales generated by Big Five-guided scaffolds. Chen et al. (2025c); Pal et al. (2025) leveraged Reddit-based journal entries to model Big Five traits in large-scale dialogue datasets. Big Five-aligned agents also improve on text based games (Lim et al., 2025), decision making under risk (Hartley et al., 2025) and code generation (Guo et al., 2025). Other efforts improve persona consistency without referencing explicit psychological theory (Takayama et al., 2025; Wu et al., 2025b). Similarly, personality is used to improve truthfulness, consistency, and context-aware generation, as further detailed in Appendix LABEL:appendix:persona_dialogue. These approaches support personality alignment but lack grounding in deeper psychological theory.
5.3 Collaborative, Multi-Agent Frameworks
Beyond task-specific capabilities, the surge in multi-agent LLM frameworks reflects a growing emphasis on collaborative decision-making, where modeling social dynamics is crucial. Social and personality psychology theories offer insights to design agent interaction, negotiation, and consensus, guiding more socially intelligent LLM systems.
@@ 5.3 Collaborative, Multi-Agent Frameworks
Persuasion models (Petty and Cacioppo, 2012) illustrate how central/peripheral routes shape attitudes in collaborative settings. Leveraging this, Gollapalli and Ng (2025) merges persuasive dialog acts with RL, Modzelewski et al. (2025) infuses persuasion knowledge into CoT, Furumai et al. (2024) combines LLM strategies and retrieval, Jin et al. (2024a); Qin et al. (2024) emphasize credibility-aware generation, and Hwang et al. (2025); Tan et al. (2025); Zeng et al. (2024a) uncovers LLMs’ vulnerabilities. Multi-agent research simulates personality-driven (Huang and Hadfi, 2024; Hu et al., 2025c; Liu et al., 2025b) and cultural-driven negotiation (Aghazadeh and Kovashka, 2025; Ma et al., 2025b), boosts truthfulness via structured debates (Khan et al., 2024), and curates argument-strength datasets (Rescala et al., 2024).
@@
ToM complements social influence by enabling agents to grasp others’ mental states. Some integrate perspective-taking (Lică et al., 2025), belief tracking (Sclar et al., 2023) and coordination (Wang et al., 2022; Sclar et al., 2022), while others refine ToM via task decomposition and recursive simulation (Jung et al., 2024; Sarangi et al., 2025; Wilf et al., 2024).
@@
Recent work on persona-driven LLM agents focuses on simulating diverse perspectives, persona alignment, and socially intelligent interactions. Han et al. (2024) introduces Big Five-based extraversion, Castricato et al. (2025) presents 1,586 synthetic personas, and Wu et al. (2025a) releases a benchmark with 40K multi-turn dialogues. Agents also model opinion dynamics (Wang et al., 2025a) and evaluate social intelligence (Chen et al., 2024b), with RoleLLM (Wang et al., 2024b), Character100 (Wang et al., 2024d), and persona-aware graph transformers (Mahajan and Shaikh, 2024) further supporting multi-party simulations. Lastly, Kumarage et al. (2025) and Chen et al. (2026) simulate social engineering attacks with LLM and Audio LM agents of varied traits, highlighting how psychological traits shape user vulnerability and the need for more robust defenses against personalized manipulation.
@@ 6.1 How Does Current LLM Development Harness Psychological Theories?
From the review from Section 1-5, we observe psychological theories have been incorporated into LLM development in stage-specific ways, with uneven coverage across theoretical domains. Figure 1 maps this integration across stages.
In early stages, preprocessing and pretraining, developmental psychology is often referenced. Its emphasis on staged knowledge acquisition aligns with curriculum learning and progressive data exposure, mirroring human developmental trajectories. In post-training, especially RLHF, behavioral psychology ideas are most prominent. Conditioning, reinforcement schedules, and reward design are commonly used to guide model alignment with human preferences. In evaluation and application, theories from social psychology, personality psychology and psycholinguistics are commonly cited, reflecting a focus on interaction patterns, user modeling, and linguistic variation – areas traditionally explored within these sub-fields. Their prominence in later stages aligns with their emphasis on human-centered communication. Cognitive psychology appears across all stages, particularly in modeling internal mechanisms such as reasoning, memory, and attention. Its breadth makes it a foundational influence, though less visible in agentic interaction settings.
The observed unevenness in integration reflects, perhaps a gap, but more probably a functional alignment – some domains are naturally better suited for certain stages of LLM development. Meanwhile, these trends expose under-explored opportunities, motivating the RQs that follow.
@@ 6.2 What Untapped Psychological Insights Could Advance LLM Development?
Although psychological theory is increasingly applied in LLM research, its use remains simplified and uneven. As shown in Tables LABEL:tab:psych_theories_dev_cog_beh, LABEL:tab:psych_theories_soc, and LABEL:tab:psych_theories_per_ling, many theories are under-utilized despite their potential to improve model behavior and interpretability. Below, we outline theories in four key areas that deserve greater attention in future LLM research.
Social psychology remains underutilized in areas like group dynamics and self and identity, limiting personalization, adaptability, and inclusivity. Prompting LLMs to adopt specific social identities can reduce bias (Dong et al., 2024a) and mirror human-like ingroup favoritism (Hu et al., 2025b). Incorporating social identity frameworks could enhance user alignment in identity-sensitive contexts (Chen et al., 2020). Likewise, while bias detection is common, classic social influence theories (e.g., conformity, obedience) and attitude change theories (e.g., balance theory, cognitive dissonance) are rarely applied to interaction dynamics or bias mitigation, despite their relevance to ethical and socially adaptive behavior. Additionally, malicious actors leveraging social influence can severely undermine trust in digital spaces (Ai et al., 2024b; Liu et al., 2025c; Zeng et al., 2024a), highlighting the potential of constructs like inoculation theory to proactively guard against manipulative strategies.
Behavioral psychology inspires RLHF, yet key concepts like partial reinforcement, which improves behavior persistence (Ferster, 1966; Jensen, 1961), and shaping, which supports gradual learning through successive approximations (Love et al., 2009), are overlooked. Current RLHF relies on uniform rewards, yet behavioral theory warns that flawed rewards can lead to reward hacking. Adding reward variability may reduce premature convergence and improve alignment with human intent (Amodei et al., 2016; Dayan and Daw, 2008).
@@
Cognitive psychology remains underused, particularly Schema Theory, which holds that humans store knowledge as schemas formed through repeated experience (Anderson and Pearson, 1984), guiding inference, memory, and learning. Recent work explores schema-inspired methods for compressing user histories and modeling knowledge activation cycles (Panagoulias et al., 2024; Xia et al., 2024). Further integration may improve long-term context handling and generalization.